OpenStudy (anonymous):

integrate (1/(x^2-3x))dx

OpenStudy (anonymous):

this is an inverse trig function

OpenStudy (anonymous):

A/ x + B/ ( x +3) = A ( x+3) + Bx -> A = -1/3, B = 1/3 => (-1/3) ln |x| + (1/3) ln | x -3| + C

OpenStudy (anonymous):

that is correct but only for restricted values of x

OpenStudy (anonymous):

Just apply what I learned :)